OpenStudy (anonymous):

solve graphically x+y=-1 and y-2x=-4..... please show work for solving equations

OpenStudy (anonymous):

\[x+y=-1\]\[ y-2x=-4\] \[x=-1-y\]\[x=2+\frac{1}{2}y\] \[-1-y=2+\frac{1}{2}y\]\[-3=\frac{3}{2}y\]\[y=-2\] \[x=-1-(-2)\]\[x=1\] \[(x,y)=(1,-2)\]